OpenStudy (anonymous):

What is the value of the product (3 – 2i)(3 + 2i)?

OpenStudy (anonymous):

9 - 4i^2 = 9 + 4 = 13 Keep in mind that i^2 = -1

OpenStudy (anonymous):

thanks